Find the focal length of a convex mirror whose radius of curvature is 32 cm.

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Radius of curvature R = 32 cm
R = 2f
f = R/2 = 32/2
= 16

Thanks